OpenStudy (anonymous):

divide. rationalize all denominators. 3+square root 18/ 1+square root 8. teacher put on the equation multiply the denominators conjugate with numerators and the denominator. how do i do this?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

i thought you did that but i wasnt sure

OpenStudy (anonymous):

\[ \frac {3+\sqrt{18}} {1+\sqrt 8}=\frac {3+\sqrt{18}} {1+\sqrt 8}\frac {1-\sqrt 8}{1-\sqrt 8} \]

OpenStudy (anonymous):

There was a misprint in the older post. Look at the new one

OpenStudy (anonymous):

yeah i was about to ask why it would be like that.

OpenStudy (anonymous):

\[ (a-b)(a+b)=a^2-b^2 \]

OpenStudy (anonymous):

ok now im confused. what do we plug into that equation

OpenStudy (anonymous):

and how do we multiply the 3+square root 18 by 1- square root 8

OpenStudy (anonymous):

\[ \frac {3+\sqrt{18}} {1+\sqrt 8}=\frac {3+\sqrt{18}} {1+\sqrt 8}\frac {1-\sqrt 8}{1-\sqrt 8}= \frac{(3+ \sqrt{18})(1-\sqrt 8)}{1^2 - (\sqrt 8)^2}=\\ \frac{(3+ \sqrt{18})(1-\sqrt 8)}{-7}\\ \]

OpenStudy (anonymous):

Expand the numerator and you are done

OpenStudy (anonymous):

so would we use foil to expand the numerator

OpenStudy (anonymous):

You see now, you do not have any radical in the denominator

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

give me one minute to work it out and if it is wrong can you help me

OpenStudy (anonymous):

um i think i got the first part but the last im not sure. 3+3square root 8+1square root 18- square root 7.07

OpenStudy (anonymous):

\[3+3\sqrt{8}+1\sqrt{18}-\sqrt{7.07}\]

OpenStudy (anonymous):

is it correct or i think the last part is wrong

OpenStudy (anonymous):

Leave the numeratro in terms of sqrt

OpenStudy (anonymous):

i lost internet sorry its taking so long my internet is being very funny today

OpenStudy (anonymous):

\[ \left(\sqrt{18}+2\right) \left(1-\sqrt{8}\right)=-10-\sqrt{2} \]

OpenStudy (anonymous):

ok so we factor it out as far down as it will go. and that is are answer

OpenStudy (anonymous):

\[ \frac{ -10-\sqrt 2}{-7}=\frac{ 10+\sqrt 2}{7} \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok is that our answer.

OpenStudy (anonymous):

I used the fact that \[ \sqrt{18} = 3 \sqrt 2\\ \sqrt{8} = 2 \sqrt 2 \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok so it is \[10+\sqrt{2}\div 7\]

OpenStudy (anonymous):

that is our answer correct

OpenStudy (anonymous):

@eliassaab can you help with my new problem i have posted